# The gauge symmetry - unbroken and broken - in the QCD ground state

### v2

@@ 3.1 Transverse relations for the full gluon self-energy
However, by this mechanism the QD constants due to the quark degrees of freedom only can be excluded from the QCD theory. What is about the exclusion of the QD constants from its YM sector? QCD being much more complicated quantum field gauge theory than QED, has the two transverse relations (3.5), second of which shows the gluon degrees of freedom explicitly. By solution of these relations, we mean that at least one of them has to be necessarily satisfied. Finally, this will lead to the removal of the corresponding QD constants from the YM theory as well as to fix the function $\xi=f(q^{2};\xi_{0})$, see exact derivations in sections 4, 5 and 6. This strategy is definitely accessible with conventional means of the gauge theories to make them renormalizable, and QCD is a such theory 2 , indeed. So that whether these transverse relations will be satisfied/saturated (i.e., equal zero) or not and further include them into the proper subtraction scheme requires much more careful investigation in QCD. Obviously, there exist only the two different requirements in order to satisfy at least one of the transverse relations (3.5). However, it is time now to introduce the proper subtraction scheme.
4 The proper subtraction scheme
The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense. It can be achieved by introducing the proper subtraction scheme in order to separate them from the PT logarithmic divergences. The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters $\tilde{\lambda}$, $\alpha$.
Within our approach nothing will depend on how exactly the regulating parameters have been introduced. Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order divergences 2 . They have to disappear from the theory after the PT and the non-perturbative (NP) renormalization programs will be performed.
Let us now introduce the subtractions for the quark and gluon contributions to the full gluon self-energy as follows:
$\Pi^{s(q)}_{\rho\sigma}(q)$ $=$ $\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$
$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)=\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$ (4.1)
@@ 4 The proper subtraction scheme
Let us remind that all the QD constants, shown up in eqs. (4.1)-(4.2), as well as the tadpole term $\Delta^{2}_{t}(D)$ itself, are independent from each other and are defined by the skeleton loop integrals, which have been already regularized from above and below. The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ 2 with $\mu^{2}\rightarrow 0$ final limit (its dimensionless counterpart is $\alpha$ mentioned above). However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example see appendix A). Then it has to be replaced by the subtraction at $q^{2}_{i}=-M^{2}_{i}$ with $M^{2}_{i}\rightarrow 0$ final limit as well, and subindex $i$ determines the number of such internal gluons. For the any full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) 26 .
Also it is necessary to emphasize that the subtractions (4.1) have not been introduced by hand. They are equivalent to add zero to the corresponding identity. For example $\Pi^{q}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)+\Pi^{q}_{\rho\sigma}(0)$, so that the initial quark contribution to the gluon self-energy is not changed. It makes it possible to present the initial quark contribution $\Pi^{q}_{\rho\sigma}(q)=\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{q}_{\rho\sigma}(0)$ as a sum of the two terms, one of which shows up all the corresponding QD but regularized constants. The first term is a regular function of the external gluon momentum $q$, by definition, and it may be only logarithmically divergent at large $q$. The same is true for the initial gluon contribution as well. Such an exact separation will be very useful for the renormalization programs of any kind to be performed for the single full gluon propagator, and we consider this as an advantage of the formalism developed here. For the interacting full gluon propagators these independent terms will interact with each other, but such a separation will simplify the corresponding renormalization programs, anyway.
The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.1), are
@@
respectively. In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the subtractions (4.1).
Substituting all these decompositions (4.3)-(4.4) into the subtractions (4.1), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(q)}_{t}(q^{2})=\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},\quad\Pi^{s(q)}_{l}(q^{2})=\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.5)
@@
$\Pi^{s(g)}_{t}(q^{2})=\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},\quad\Pi^{s(g)}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.6)
@@
Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.1) above, as it has to be. Also these relations will be true if expressed in terms of the above-mentioned subtraction point $q^{2}=-\mu^{2}$.
Let us now make an important observation in advance. If the initial contributions of the quarks and gluons degrees of freedom to the full gluon self-energy, which are present by the second lines in the relations (4.3) and (4.4), respectively, are not transverse, then one cannot remove the corresponding QD constants $\Delta^{2}_{q}$ and $\Delta^{2}_{g}(D)$ from the theory. On the contrary, if their are transverse, i.e., $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$, then from the relations (4.5) and (4.6) it follows that one has to put the corresponding QD constants $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus removing them from the theory, since the corresponding subtracted invariant functions cannot have the pole-type singularities, by definition. This observation explicitly shows the necessity to satisfy one of the transverse relations (3.5) from the general point of view.
@@
An interesting observation follows from the relations (4.5) and (4.6) as well, namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$. So that the corresponding sums do not depend on the corresponding QD but regularized constants, which once more underlines the general character of our subtraction scheme. Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.1).
5 Gauge symmetry is unbroken
In order to satisfy the first transverse relation (3.5) one needs to put $\xi=f(\xi_{0})=\xi_{0}$, and thus it becomes transverse $q_{\rho}q_{\sigma}\Pi_{\rho\sigma}(q;D)=0$, indeed. Then from the second of the relations (3.5) it follows that
$q_{\rho}q_{\sigma}[\Pi^{q}_{\rho\sigma}(q)+\Pi^{g}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}_{t}(D),$ (5.1)
and it is not transverse, as expected. Substituting further the subtractions (4.1) into the relation (5.1), and doing some algebra, one obtains
$q_{\rho}q_{\sigma}[\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{s(g)}_{\rho\sigma}(q;D)]=-q^{2}\Delta^{2}(D),$ (5.2)
@@ 5 Gauge symmetry is unbroken
$\Delta^{2}(D)=\Delta^{2}_{q}+\Delta^{2}_{g}(D)+\Delta^{2}_{t}(D),$ (5.3)
is the sum of all the QD but regularized constants, which may appear in the theory.The QD but regularized constant $\Delta^{2}_{g}(D)$ itself is defined in eq. (4.2).
Substituting further the first tensor decompositions of the relations (4.3) and (4.4) into the transverse relation (5.2), and doing some algebra, one obtains for their longitudinal invariant functions the following relation
$\Pi^{s(q)}_{l}(q^{2})+\Pi^{s(g)}_{l}(q^{2})={\Delta^{2}(D)\over q^{2}},$ (5.4)
which is impossible, since these invariant functions cannot have the pole-type singularities, by definition, as underlined just above. So that the QD but regularized constant $\Delta^{2}(D)$ should be discarded from the theory, i.e., put formally zero $\Delta^{2}(D)=0$ on this mathematical basis. Then $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2})=0$ as well, since they are independent from each other.
From now on in this section, we replace $D$ by $D^{PT}$ (for the explanation see the text below). The definitions (4.2) and (5.3) require that this system of the relations for the QD but regularized constants now looks like
$\Delta^{2}_{q}=\Delta^{2}_{g}(D^{PT})=\Delta^{2}_{t}(D^{PT})=0,$ (5.5)

### v3

@@ 3.2 Particular solution
but these equalities should be put by hand in this case, i.e., they are the prescriptions rather than the exact mathematical results. We will call it as the particular solution to QCD. Let us note in advance that these conditions determine the so-called conventional or, equivalently, conventional PT QCD (though the coupling constant remains always strong apart from the asymptotic free (AF) regime [2, 38, 39]). Precisely, this solution to QCD has been studied the last five decades. In this solution the gauge symmetries of the QCD Lagrangian and its ground state coincide, like in QED (see section 6 as well).
4 Proper subtraction scheme
QCD being much more complicated quantum field gauge theory than QED, has the two satisfied transverse relations (3.14), and so requires much more careful investigation in QCD. They have to be included in a self-consistent way into the proper subtraction scheme. Finally this will lead to the removal of the corresponding QD constants from the theory. How they fixed the gauge-fixing function approach to QCD in the most general and unique ways has been just described here.
The first step in the renormalization program of any gauge theory is the removal of the quadratic UV divergences in order to make the corresponding theory renormalizable in the PT sense. It can be achieved by introducing the proper subtraction scheme in order to separate them from the PT logarithmic divergences. The preliminary step in the regularization program has been already done by introducing the corresponding regulating parameters, mentioned above. Within our approach nothing will depend on how exactly the regulating parameters have been introduced. Even the hard cut-off procedures can be used, since we are mainly interested in the regulating of leading order UV divergences [2]. They have to disappear from the theory after the PT and the non-perturbative (NP) renormalization programs will be performed. Let us now introduce the subtractions for the quark and gluon contributions to the full gluon self-energy as follows:
$\Pi^{s(q)}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)=\Pi^{q}_{\rho\sigma}(q)-\delta_{\rho\sigma}\Delta^{2}_{q},$ (4.18)
$\Pi^{s(g)}_{\rho\sigma}(q;D)$ $=$ $\Pi^{g}_{\rho\sigma}(q;D)-\Pi^{g}_{\rho\sigma}(0;D)$ (4.19)
$=$ $\Pi^{g}_{\rho\sigma}(q;D)-\delta_{\rho\sigma}\Delta^{2}_{g}(D),$
@@ 4 Proper subtraction scheme
Let us remind that all the QD constants, shown up in eqs. (4.18)-(4.19), as well as the tadpole term $\Delta^{2}_{t}(D)$ itself, are independent from each other and are defined by the skeleton loop integrals, which have been already regularized from above and below. The subtraction at zero is to be understood in a such way that we subtract at the external gluon momentum $q^{2}=-\mu^{2}$ [2] with $\mu^{2}\rightarrow 0$ final limit. However, due to the self-interaction of the massless gluon modes, some of these constants (apart from the tadpole term) may depend on the internal zero gluon momentum (if it is connected to any closed loop, for example such as which can appear in the iteration solution of the gluon SD equation [6]). Then it has to be replaced by the subtraction at $q^{2}_{i}=-M^{2}_{i}$ with $M^{2}_{i}\rightarrow 0$ final limit as well, and subindex $i$ determines the number of such internal gluons. For the any single full gluon propagator the subindex $i$ can be treated as the number of the necessary subtractions made in it in accordance with the rules of the theory of distributions (generalized functions) [26].
Also it is necessary to emphasize that the subtractions (4.17)-(4.18) have not been introduced by hand. They are equivalent to add zero to the corresponding identity. For example $\Pi^{q}_{\rho\sigma}(q)=\Pi^{q}_{\rho\sigma}(q)-\Pi^{q}_{\rho\sigma}(0)+\Pi^{q}_{\rho\sigma}(0)$, so that the initial quark contribution to the gluon self-energy is not changed. It makes it possible to present the initial quark contribution $\Pi^{q}_{\rho\sigma}(q)=\Pi^{s(q)}_{\rho\sigma}(q)+\Pi^{q}_{\rho\sigma}(0)$ as a sum of the two terms, one of which shows up all the corresponding QD but regularized constants. The first term is a regular function of the external gluon momentum $q$, by definition, and it may be only logarithmically divergent at large $q$. In other words, we separate the NP QD constants from the PT logarithmically divergent terms in a gauge-invariant way, that is why our subtraction scheme is called as a proper one. The same is true for the initial gluon contribution as well. Such an exact separation will be very useful for the renormalization programs of any kind to be performed for the single full gluon propagator, and we consider this as an advantage of the formalism developed here. For the interacting full gluon propagators these independent terms will interact with each other, but such a separation will simplify the corresponding renormalization programs, anyway.
The independent tensor decompositions of the quark and gluon degrees of freedom, which appear in the subtraction relations (4.18)-(4.19), are
@@
respectively. In all the quantities above and below the dependence on $D$ is omitted, for simplicity, and will be restored when necessary. Here and everywhere below all the invariant functions are dimensionless ones of their argument $q^{2}$: otherwise they remain arbitrary. However, all the invariant functions $\Pi^{s(q)}_{t}(q^{2})$, $\Pi^{s(q)}_{l}(q^{2})$ as well as $\Pi^{s(g)}_{t}(q^{2})$, $\Pi^{s(g)}_{l}(q^{2})$ cannot have the power-type singularities (or, equivalently, the pole-type or the massless ones) at small $q^{2}$, by definitions, as it follows from the initial subtractions (4.18)-(4.19).
Substituting all these decompositions (4.21)-(4.22) into the subtractions (4.18)-(4.19), and doing some tensor algebra derivations, one finally obtains
$\Pi^{s(q)}_{t}(q^{2})$ $=$ $\Pi^{q}_{t}(q^{2})-{\Delta^{2}_{q}\over q^{2}},$
$\Pi^{s(q)}_{l}(q^{2})$ $=$ $\Pi^{q}_{l}(q^{2})+{\Delta^{2}_{q}\over q^{2}},$ (4.23)
@@
$\Pi^{s(g)}_{t}(q^{2})$ $=$ $\Pi^{g}_{t}(q^{2})-{\Delta^{2}_{g}(D)\over q^{2}},$
$\Pi^{s(g)}_{l}(q^{2})$ $=$ $\Pi^{g}_{l}(q^{2})+{\Delta^{2}_{g}(D)\over q^{2}}.$ (4.24)
@@
Using these relations it is easy to show that for the regularized (i.e., finite) quark and gluon contributions to the initial gluon self-energy their values at zero always are $\Pi^{q}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{q}$ and $\Pi^{g}_{\rho\sigma}(0;D)=\delta_{\rho\sigma}\Delta^{2}_{g}(D)$, indeed, i.e, they are general ones, and have been already used in the subtraction relations (4.18)-(4.19) above, as it has to be. Also these relations will be true if expressed in terms of the above-mentioned subtraction point $q^{2}=-\mu^{2}$.
@@
An interesting observation follows from the relations (4.23) and (4.24), namely $\Pi^{q}_{t}(q^{2})+\Pi^{q}_{l}(q^{2})=\Pi^{s(q)}_{t}(q^{2})+\Pi^{s(q)}_{l}(q^{2})$ and $\Pi^{g}_{t}(q^{2})+\Pi^{g}_{l}(q^{2})=\Pi^{s(g)}_{t}(q^{2})+\Pi^{s(g)}_{l}(q^{2})$. So that the corresponding sums do not depend on the corresponding QD but regularized constants, which once more underlines the general character of our subtraction scheme. Also, the sum of the subtracted invariant functions is never zero, though the subtracted contributions to the full gluon self-energy are zero at $q^{2}=0$, by definitions, see relations (4.18)-(4.19).
As we already know, the quark and gluon contributions to the full gluon self-energy are transverse $q_{\rho}q_{\sigma}\Pi^{q}_{\rho\sigma}(q)=q_{\rho}q_{\sigma}\Pi^{g}_{\rho\sigma}(q;D)=0$, see the relations (3.14). Then from the second of the relations (4.21) and (4.22) it follows that $\Pi^{q}_{l}(q^{2})=\Pi^{g}_{l}(q^{2})=0$ as well. So that the second relations in (4.23) and (4.24) will be reduced to
where we restored the dependence on $D$ in the second relation. However, these relations are impossible, since the corresponding invariant functions cannot have the pole-type singularities, by definitions, as explained above. One has to put these constants to zero on this general mathematical basis, i.e., $\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0$, and thus $\Pi^{s(q)}_{l}(q^{2})=\Pi^{s(g)}_{l}(q^{2};D)=0$ as well. All our results for the QD constants can be now summarized as follows:
$\Delta^{2}_{q}=\Delta^{2}_{g}(D)=0,\quad\Delta^{2}_{t}(D)\neq 0,$ (4.26)
and hence $\Delta^{2}_{gh}=\Delta^{2}_{1}(D^{2})=\Delta^{2}_{2}(D^{4})=\Delta^{2}_{2^{\prime}}(D^{3})=0$, as well because of the relation (4.20). The system of the relations (4.26) is not only a general but it is a unique one as well. They are not prescriptions, since based on a rigorous tensor algebra derivation rules, i.e., they are exact mathematical results. The characteristic feature of this system is that the tadpole term $\Delta_{t}^{2}(D)$ remains intact. The mathematical reason is that for the constant tadpole term such kind of the non-trivial relations (4.18)-(4.19) does not exist (constant minus the same constant is always zero). The essential dynamical source of its survival in the theory is that only the tadpole term can generate a mass squared scale parameter [6]. All the other quark and gluon terms, contributing to the full gluon self-energy, cannot do this, and thus none of their subtracted counterparts with their quark and gluon constants.
All these constants, defined by the corresponding skeleton loop integrals, are QD at the upper limit. They have to be removed/disregarded from the theory on the general mathematical basis, i.e., put zero, as described above, apart from the tadpole term (2.4). In comparison with it, all the other QD quark and gluon constants will be called the tadpole-like/type terms. The general question arises now, namely how to understand these exact equalities to zero in the relations (4.26)? These equalities mean that any tadpole-like term which may appear in the theory by any possible way has to be discarded/disregarded in the theory, i.e, put zero, independently from any other tadpole-like terms. In this way such sums will be always zero, indeed. Such constants may appear even not as the result of the subtractions, but, for example as the result of the NL iteration procedure for the full gluon propagator [6] (see appendix A as well). In this case the tadpole-like terms may be even multiplied by some regularized functions. Whatever their origins would have been in the full gluon propagator, all of them belong to the infinite manifold of the relations (4.26), and thus should be always removed from the theory.
Collecting our results obtained above for the quark and gluon contributions to the gluon self-energy, one obtains
$\Pi^{q}_{\rho\sigma}(q)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(q)}_{t}(q^{2}),$
$\Pi^{g}_{\rho\sigma}(q;D)$ $=$ $T_{\rho\sigma}(q)q^{2}\Pi^{s(g)}_{t}(q^{2};D),$ (4.27)
since they become transverse and coincide with their subtracted counterparts. Reminding also that the invariant functions present in the previous relations are regular functions at small $q^{2}$, i.e., they have no pole-type singularities, and may be only logarithmically divergent at large $q^{2}$. This becomes possible only due to the satisfied transverse relations (3.14) for the quark and gluon degrees of freedom. Just this decreases the quadratic UV divergences of the corresponding skeleton loop integrals to a logarithmic ones, as it has been described in detail in this section.
5 General solution to QCD
Substituting the sum of the terms (4.27) into the eq. (3.12), on account of the relation $\delta_{\rho\sigma}=T_{\rho\sigma}(q)+L_{\rho\sigma}(q)$ and further into the gluon SD eq. (2.5), one finally gets
$D_{\mu\nu}(q)=D^{0}_{\mu\nu}(q)+D^{T}_{\mu\nu}(q;D)+D^{L}_{\mu\nu}(q;D),$ (5.28)
@@ 5 General solution to QCD
$D^{T}_{\mu\nu}(q;D)=D^{0}_{\mu\rho}(q)iT_{\rho\sigma}(q)[q^{2}\Pi(q^{2};D)+\Delta^{2}_{t}(D)]D_{\sigma\nu}(q),$ (5.29)